# A theoretical analysis on the inversion of matrices via Neural Networks designed with Strassen algorithm

### v3

@@ 2.1 Linear Algebra
In this paper, we denote the spaces of matrices of $n_{1}$ rows and $n_{2}$ columns by $\mathbb{R}^{n_{1}\times n_{2}}$. Given $A\in\mathbb{R}^{n_{1}\times n_{2}}$, $i\in\left\{1,\dots,n_{1}\right\}$ and $j\in\left\{1,\dots,n_{2}\right\}$, $A_{i,j}$ denotes the entry in its $i$-th row and $j$-th column. Moreover, given $c_{1},c_{2}\in\left\{1,\dots,n_{1}\right\}$ such that $c_{1}\leq c_{2}$ and $c_{3},c_{4}\in\left\{1,\dots,n_{2}\right\}$ such that $c_{3}\leq c_{4}$ we denote by $A_{c_{1}:c_{2},c_{3}:c_{4}}$ its submatrix formed by the rows from $c_{1}$ to $c_{2}$ (including $c_{1}$ and $c_{2}$), and the columns from $c_{3}$ to $c_{4}$ (including $c_{3}$ and $c_{4}$). We also denote by $\operatorname{I}_{n}\in\mathbb{R}^{n\times n}$ the identity matrix whose entries $(\operatorname{I}_{n})_{i,j}=\delta_{i,j}$ for all $i,j\in\left\{1,\dots,n\right\}$.
@@
Finally, we denote by $\left\lvert\;\cdot\;\right\rvert_{p}$ the usual $p$ norm for vectors and matrices (seen as a flat vectors) and by $\left\lVert\;\cdot\;\right\rVert_{p}$ the matrix norm induced by $\left\lvert\;\cdot\;\right\rvert_{p}$.
In particular, we need the following estimates between norms:
@@ 2.2 Strassen Algorithm for Square Matrices
In this section, we recall the Strassen algorithm, which was proposed in [strassen1969gaussian]. First, we define the Strassen algorithm in matrices belonging to $\mathbb{R}^{2^{k}\times 2^{k}}$. Then, we define it in matrices in $\mathbb{R}^{n\times n}$ for arbitrary values $n\in\mathbb{N}$.
@@ 2.3 Matrix Neural Networks
We now define the notions of Matrix Neural Network, concatenation, and parallelization. This is based on the definition introduced in [petersen2018optimal] and [gribonval2022approximation], but adapted for Neural Networks whose entries are matrices. We do this for the sake of simplifying the notation.
@@
Multilayer Perceptron (if we consider them defined as in [petersen2018optimal]) have the same structure as a Matrix Neural Networks: they are a sequence of triples of matrices, vectors, and activation functions. Their realization, number of layers, and number of weights are defined equivalently for these Neural Networks. In particular, any Multilayer Perceptron can be viewed as a Matrix Neural Network by considering $N^{\ell,2}=1$ for all $\ell=1,\dots,N$. Conversely, using vectorization procedures, since the matrix space $\mathbb{R}^{n\times m}$ is isomorphic to the vector space $\mathbb{R}^{nm}$, all the Matrix Neural Networks can be implemented as classical Neural Networks (with the same length, number of weights, and final error). Therefore, we can view both notions as equivalent.
We can directly extend for a $\varrho$-MNN $\Phi$ the definitions given in [kutyniok2022theoretical] of its input and output dimensions $\operatorname{dim_{in}}(\Phi),\operatorname{dim_{out}}(\Phi)$, number of layers $L(\Phi)$, number of weights at the layer $\ell$ $M_{\ell}(\Phi)$, number of weights $M(\Phi)$ and realization $R(\Phi)$. We omit their formal definitions for the sake of brevity as they are standard notions.
For more than one $\varrho$-MNN $\Phi^{1},\dots,\Phi^{k}$, we can again extend the ideas of their sparse concatenation $\Phi^{1}\odot\Phi^{2}$ and their parallelization $\mathcal{P}(\Phi^{1},\dots,\Phi^{k})$ as defined in, for example, [petersen2018optimal] and [kutyniok2022theoretical]. In these papers, they focused in the activation function $\operatorname{ReLU}$ and these definitions are written with that in mind, but, allowing components of the activation functions to be the identity solves this technicalities. The whole point of introducing these concepts is that they satisfy the following:
@@
In order to implement the Strassen algorithm with $\varrho$-MNN, with $\varrho$ fixed, we consider that we may multiply two numbers using Neural Networks. In particular, we suppose the following.
@@
This is easily satisfied by the multiple versions of the Universal Approximation Theorem, for example, that in [cybenko1989approximation]. Actually, Hypothesis 1 is equivalent to these results: the product approximation unlocks the polynomial approximation and therefore, the approximation of continuous functions.
@@
For the particular case $\varrho=\operatorname{ReLU}=t\mapsto\max\left\{t,0\right\}$, $\Pi_{\varepsilon,K}^{\operatorname{ReLU}}$ is constructed in [elbrachter2022dnn], which, by a minor modification using the equality
@@
Remark 2.6.
3 Main Results
3.1 Matrix Multiplication Approximation
@@
$\Phi^{\operatorname{PAR}_{k,\varrho}}_{\varepsilon,K}=\mathcal{P}\left(\Phi^{\operatorname{STR}_{2^{k-1},\varrho}}_{\frac{\varepsilon}{4},2K},\Phi^{\operatorname{STR}_{2^{k-1},\varrho}}_{\frac{\varepsilon}{4},2K},\Phi^{\operatorname{STR}_{2^{k-1},\varrho}}_{\frac{\varepsilon}{4},2K},\right.\\ \left.\Phi^{\operatorname{STR}_{2^{k-1},\varrho}}_{\frac{\varepsilon}{4},2K},\Phi^{\operatorname{STR}_{2^{k-1},\varrho}}_{\frac{\varepsilon}{4},2K},\Phi^{\operatorname{STR}_{2^{k-1},\varrho}}_{\frac{\varepsilon}{4},2K},\Phi^{\operatorname{STR}_{2^{k-1},\varrho}}_{\frac{\varepsilon}{4},2K}\right).$
@@
For a detailed proof, see Appendix A.
Let us now extend this results to matrices of any shape. Consider $m,n,p\in\mathbb{N}$. In order to multiply a matrix $A\in\mathbb{R}^{m\times n}$ by another matrix $B\in\mathbb{R}^{n\times p}$ using the Strassen algorithm implemented with Neural Networks, we are going to prolong the matrices to size $2^{k}\times 2^{k}$, for:
@@
This result is a direct consequence of 3.1 and the definition of $\Phi^{\operatorname{STR}}$ and $\Phi^{\operatorname{EXT}}$.
@@
Compared to the result presented in [kutyniok2022theoretical, Proposition 3.7], we reduce the dimension dependency of order 3 to $\log_{2}7$, as expected by applying Strassen’s algorithm.
@@
When $n=2^{k}$, $R\left(\Phi^{\operatorname{SHR}_{n}}_{n,n}\right)$ and $R\left(\Phi_{n}^{\operatorname{EXT}^{*}}\right)$ are identity tensors, which is why we consider (3.14) an extension. In that case, there is only a small difference in its length and number of weights, which is irrelevant for studying their asymptotic values.
@@
3.2 Matrix Inversion Approximation
@@
The Neumann series can be used to approximate the inverse of any matrix since a matrix $A$ is invertible if and only if there is a (invertible) matrix $P$ for which $\left\lVert\operatorname{I}-AP\right\rVert_{2}<1$. In that case,
@@
The particular case when $P=\alpha\operatorname{I}$ is of high interest for symmetric positive definite matrices, as $\alpha$ can be estimated fairly easily.
@@
The proof can be found in Appendix B. The idea of the proof is first construct a MNN that approximates the powers $A^{2^{k}}$, then another responsible of the product (3.19) and finally, using the Neumann series for an appropriate number of terms ($2^{N(\frac{\varepsilon}{2\alpha},\delta)+1}$ concretely), we get the desired MNN.
@@
As stated in Remark 3.6, we could construct a $\varrho$-MNN that approximates any matrix $A\in\mathbb{R}^{d\times d}$ such that
$\left\lVert\operatorname{I}_{d}-AP\right\rVert\leq\delta<1,$
where is $P\in\mathbb{R}^{d\times d}$ is a fixed invertible matrix. But for simplicity of the statement and proof, we stick to the case $P=\alpha\operatorname{I}_{d}$.
@@
For the particular case $\varrho=\operatorname{ReLU}$ and $\alpha=1$, estimates (3.22) and (3.23) can be written as
@@
This result presents various improvements with respect to prior work, concretely [kutyniok2022theoretical, Theorem 3.8]. Although the number of layers is asymptotically worse, the upper bound on the number of weights on (3.26) is an improvement over theirs. The exponent on the dimension of the input matrices is reduced from 3 to $\log_{2}7$ and the factor is $\log_{2}m(\varepsilon,\delta)$ instead of $m(\varepsilon,\delta)\log_{2}^{2}(m(\varepsilon,\delta))$. In parentheses, even if we have $m(\varepsilon,\delta)$ instead of $\log_{2}m(\varepsilon,\delta)$, it is still an improvement if we compare the expanded expression $m(\varepsilon,\delta)\log_{2}(m(\varepsilon,\delta))$ to $m(\varepsilon,\delta)\log_{2}(m(\varepsilon,\delta))\log_{2}^{2}(m(\varepsilon,\delta))$. The first change comes from the use of the Strassen algorithm for matrix multiplication, while the second is due to the use of the identity (3.19).
@@
In [kutyniok2022theoretical], they consider parametric coercive PDEs whose discretizations, by applying the Galerkin method, results in a linear system whose respective matrices are symmetric positive definite. They prove that the inverse of all of those matrices can be approximated by a single Neural Network (with worse upper bounds on the number of parameters as explained in Remark 3.9). Then, we reach the same conclusion with a better dependency on the Galerkin dimension for approximating the parameter-to-solution map. However, other methods, like those of multilevel ideas in [JMLR:v24:23-0421], provides even better bounds for the particular equation considered where they relay in a more efficient way of solving the resulting linear system.
@@ 4 Conclusions
We have constructed two important Neural Networks approximating matrix multiplication and inversion operators. We have been able to do so without restricting ourselves to a fixed activation function. Also, we bound their number of weights and layers, and by using the Strassen algorithm we have reduced a potential cubic dependency on the dimension to a $\log_{2}7$ exponent. Rewriting the Neumann partial sum, we reduce the upper bounds on the number of parameters by reducing the number of matrix multiplication apparitions.
@@
[…]
@@
[…]
@@
Part of the improvements with respect to the paper [kutyniok2022theoretical] can be found in the Master Thesis [romera2024neuralnetworksnumericalanalysis]
@@
Step 1: The simple case $N=1$. We just have define $\Phi^{\operatorname{NEU}^{1}_{n,\varrho}}_{\varepsilon}=((\operatorname{id},\operatorname{I}_{n},\operatorname{id}))$, from where the result follows easily.
@@
We do this to have upper bounds of the approximations with the cost of the power of two as a factor.
@@
$\left\lVert\left(\operatorname{SQR}_{\varepsilon}^{i}(A)+2^{-2^{i}}\operatorname{I}_{n}\right)\operatorname{PROD}_{\varepsilon}^{i}(A)\right.\\ \left.-R\left(\Phi_{\frac{\varepsilon}{4n},1}^{\operatorname{STR}_{n,\varrho}}\right)\left(\operatorname{SQR}_{\varepsilon}^{i}(A)+2^{-2^{i}}\operatorname{I}_{n}\middle|\operatorname{PROD}_{\varepsilon}^{i}(A)\right)\right\rVert_{2}\leq\frac{\varepsilon}{4}.$
@@
We first focus on proving (3.21). We know that the identity

### v4

@@ 2.1 Linear Algebra
In this paper, we denote the spaces of matrices of $n_{1}$ rows and $n_{2}$ columns by $\mathbb{R}^{n_{1}\times n_{2}}$. Given $A\in\mathbb{R}^{n_{1}\times n_{2}}$, $i\in\left\{1,\dots,n_{1}\right\}$ and $j\in\left\{1,\dots,n_{2}\right\}$, $A_{i,j}$ denotes the entry in its $i$-th row and $j$-th column. Moreover, given $c_{1},c_{2}\in\left\{1,\dots,n_{1}\right\}$ such that $c_{1}\leq c_{2}$ and $c_{3},c_{4}\in\left\{1,\dots,n_{2}\right\}$ such that $c_{3}\leq c_{4}$ we denote by $A_{c_{1}:c_{2},c_{3}:c_{4}}$ its submatrix formed by the rows from $c_{1}$ to $c_{2}$ (including $c_{1}$ and $c_{2}$), and the columns from $c_{3}$ to $c_{4}$ (including $c_{3}$ and $c_{4}$). We also denote by $\operatorname{I}_{n}\in\mathbb{R}^{n\times n}$ the identity matrix whose entries are $(\operatorname{I}_{n})_{i,j}=\delta_{i,j}$ for all $i,j\in\left\{1,\dots,n\right\}$.
@@
Finally, we denote by $\left\lvert\;\cdot\;\right\rvert_{p}$ the usual $p$ norm for vectors and matrices (seen as vectors) and by $\left\lVert\;\cdot\;\right\rVert_{p}$ the matrix norm induced by $\left\lvert\;\cdot\;\right\rvert_{p}$.
In particular, we need the following classical estimates between norms:
@@ 2.2 Strassen Algorithm for Square Matrices
In this section, we recall the Strassen algorithm, which was proposed in [47]. First, we define the Strassen algorithm in matrices belonging to $\mathbb{R}^{2^{k}\times 2^{k}}$. Then, we define it in matrices in $\mathbb{R}^{n\times n}$ for arbitrary values $n\in\mathbb{N}$.
@@ 2.3 Matrix Neural Networks
We now define the notions of Matrix Neural Network, concatenation, and parallelization. This is based on the definition introduced in [40] and [18], but adapted for Neural Networks whose entries are matrices. We do this for the sake of simplifying the notation.
@@
Multilayer Perceptron (if we consider them defined as in [40]) have the same structure as a Matrix Neural Networks: they are a sequence of triples of matrices, vectors, and activation functions. Their realization, number of layers, and number of weights are defined equivalently for these Neural Networks. In particular, any Multilayer Perceptron can be viewed as a Matrix Neural Network by considering $N^{\ell,2}=1$ for all $\ell=1,\dots,N$. Conversely, using vectorization procedures, since the matrix space $\mathbb{R}^{n\times m}$ is isomorphic to the vector space $\mathbb{R}^{nm}$, all the Matrix Neural Networks can be implemented as classical Neural Networks (with the same length, number of weights, and final error). Therefore, we can view both notions as equivalent.
We can directly extend for a $\varrho$-MNN $\Phi$ the definitions given in [30] of its input and output dimensions $\operatorname{dim_{in}}(\Phi),\operatorname{dim_{out}}(\Phi)$, number of layers $L(\Phi)$, number of weights at the layer $\ell$ $M_{\ell}(\Phi)$, number of weights $M(\Phi)$ and realization $R(\Phi)$. We omit their formal definitions for the sake of brevity as they are standard notions.
For more than one $\varrho$-MNN $\Phi^{1},\dots,\Phi^{k}$, we can again extend the ideas of their sparse concatenation $\Phi^{1}\odot\Phi^{2}$ and their parallelization $\mathcal{P}(\Phi^{1},\dots,\Phi^{k})$ as defined in, for example, [40] and [30]. In these papers, they focused in the activation function $\operatorname{ReLU}$ and these definitions are written with that in mind, but, allowing components of the activation functions to be the identity solves this technicalities. The whole point of introducing these concepts is that they satisfy the following:
@@
In order to implement the Strassen algorithm with $\varrho$-MNN, we suppose that we may multiply two numbers using Neural Networks:
@@
This is easily satisfied by the multiple versions of the Universal Approximation Theorem, for example, that in [8]. Actually, Hypothesis 1 is equivalent to these results: the product approximation unlocks the polynomial approximation and therefore, the approximation of continuous functions.
@@
Remark 2.6.
For the particular case $\varrho=\operatorname{ReLU}=t\mapsto\max\left\{t,0\right\}$, $\Pi_{\varepsilon,K}^{\operatorname{ReLU}}$ is constructed in [12], which, by a minor modification using the equality
@@
3 Matrix Multiplication Approximation
@@
$\Phi^{\operatorname{PAR}_{k,\varrho}}_{\varepsilon,K}=\mathcal{P}\left(\Phi^{\operatorname{STR}_{2^{k-1},\varrho}}_{\frac{\varepsilon}{4},2K},\Phi^{\operatorname{STR}_{2^{k-1},\varrho}}_{\frac{\varepsilon}{4},2K},\Phi^{\operatorname{STR}_{2^{k-1},\varrho}}_{\frac{\varepsilon}{4},2K},\Phi^{\operatorname{STR}_{2^{k-1},\varrho}}_{\frac{\varepsilon}{4},2K},\Phi^{\operatorname{STR}_{2^{k-1},\varrho}}_{\frac{\varepsilon}{4},2K},\Phi^{\operatorname{STR}_{2^{k-1},\varrho}}_{\frac{\varepsilon}{4},2K},\Phi^{\operatorname{STR}_{2^{k-1},\varrho}}_{\frac{\varepsilon}{4},2K}\right).$
@@
The result is proved in Appendix A.
Let us now extend this result to matrices of any shape. Consider $m,n,p\in\mathbb{N}$. In order to multiply a matrix $A\in\mathbb{R}^{m\times n}$ by another matrix $B\in\mathbb{R}^{n\times p}$ using the Strassen algorithm implemented with Neural Networks, we are going to prolong the matrices to size $2^{k}\times 2^{k}$, for:
@@
This result is a direct consequence of Proposition 3.1 and the definition of $\Phi^{\operatorname{STR}}$ and $\Phi^{\operatorname{EXT}}$.
@@
Compared to the result presented in [30, Proposition 3.7], we reduce the dimension dependency of order 3 to $\log_{2}7$, as expected by applying Strassen’s algorithm.
@@
When $n=2^{k}$, $R\left(\Phi^{\operatorname{SHR}_{n}}_{n,n}\right)$ and $R\left(\Phi_{n}^{\operatorname{EXT}^{*}}\right)$ are identity tensors, which is why we consider (3.14) an extension. In that case, there is only a small difference in its length and number of weights, which is irrelevant for getting the asymptotic values on the number of weights and layers.
@@
4 Matrix Inversion Approximation
@@
The Neumann series can be used to approximate the inverse of any invertible matrix $A\in\mathbb{R}^{n\times n}$. If $P\in\mathbb{R}^{n\times n}$ satisfies $\left\lVert\operatorname{I}_{n}-AP\right\rVert_{2}<1$ then
@@
Conversely, if such $P$ does not exist, then $A$ cannot be invertible. The particular case where $P=\alpha\operatorname{I}$ is of high interest for real symmetric positive definite matrices, as $\alpha$ can be estimated fairly easily. Indeed, if $A$ is real symmetric, it can be diagonalized and
$\left\lVert\operatorname{I}_{n}-\alpha A\right\rVert_{2}=\max_{\lambda\in\Lambda(A)}|1-\alpha\lambda|,$
for $\Lambda(A)$ the set of eigenvalues of $A$. Real symmetric positive definite matrices appear naturally in the resolution of PDEs with the Galerkin method or finite differences.
@@
The proof can be found in Appendix B. The idea of the proof is first construct a MNN that approximates the powers $A^{2^{k}}$, then another to compute the product (4.2), and finally, using the Neumann series for an appropriate number of terms ($2^{N(\frac{\varepsilon}{2\alpha},\delta)+1}$ concretely), we get the desired MNN.
@@
As stated in Remark 4.1, we could construct a $\varrho$-MNN that approximates any matrix $A\in\mathbb{R}^{n\times n}$ such that
$\left\lVert\operatorname{I}_{n}-AP\right\rVert\leq\delta<1,$
where is $P\in\mathbb{R}^{n\times n}$ is a fixed invertible matrix. But for simplicity of the statement and proof, we stick to the case $P=\alpha\operatorname{I}_{n}$.
@@
For the particular case $\varrho=\operatorname{ReLU}$ and $\alpha=1$, estimates (4.5) and (4.6) can be written as:
@@
This result presents various improvements with respect to prior work, concretely [30, Theorem 3.8]. Although the number of layers is asymptotically worse, the upper bound on the number of weights on (4.9) is an improvement over theirs. The exponent on the dimension of the input matrices is reduced from 3 to $\log_{2}7$ and the factor is $\log_{2}m(\varepsilon,\delta)$ instead of $m(\varepsilon,\delta)\log_{2}^{2}(m(\varepsilon,\delta))$. In parentheses, even if we have $m(\varepsilon,\delta)$ instead of $\log_{2}m(\varepsilon,\delta)$, it is still an improvement if we compare the expanded expression $m(\varepsilon,\delta)\log_{2}(m(\varepsilon,\delta))$ to $m(\varepsilon,\delta)\log_{2}(m(\varepsilon,\delta))\log_{2}^{2}(m(\varepsilon,\delta))$. The first change comes from the use of the Strassen algorithm for matrix multiplication, while the second is due to the use of the identity (4.2).
@@
In [30], they consider parametric coercive PDEs whose discretizations, by applying the Galerkin method, results in a linear system whose respective matrices are symmetric positive definite. They prove that the inverse of all of those matrices can be approximated by a single Neural Network (with worse upper bounds on the number of parameters as explained in Remark 4.4). Then, we reach the same conclusion with a better dependency on the Galerkin dimension for approximating the parameter-to-solution map. However, other methods, like those of multilevel ideas in [24], provides even better bounds for the particular equation considered where they rely on a more efficient way of solving the resulting linear system.
@@ 5 Conclusions
We have constructed two Neural Networks approximating matrix multiplication and inversion operators. We have been able to do so without restricting ourselves to a fixed activation function. Also, we bound their number of weights and layers, and by using the Strassen algorithm we have reduced a potential cubic dependency on the dimension to a $\log_{2}7$ exponent. Rewriting the Neumann partial sum, we reduce the upper bounds on the number of parameters by reducing the number of matrix multiplication apparitions.
@@
$\left|\left[R\left(\Phi^{\operatorname{PAR}_{k,\varrho}}_{\varepsilon,K}\odot\Phi^{\operatorname{SPLIT}_{k}}\right)(A|B)\right]_{(i-1)\cdot 2^{k-1}+1:i\cdot 2^{k-1},1:2^{k-1}}-P_{i}\right|_{\infty}\leq\frac{\varepsilon}{4}\quad\forall i\in\left\{1,\dots,7\right\}.$ (A.1)
@@
$\left\lvert R\left(\Phi^{\operatorname{STR}_{2^{k-1},\varrho}}_{\frac{\varepsilon}{4},2K}\right)\left(A[1,1]+A[2,2]\middle|B[1,1]+B[2,2]\right)\right.\left.-(A[1,1]+A[2,2])(B[1,1]+B[2,2])\right\rvert_{\infty}\leq\frac{\varepsilon}{4},$
@@
Part of the improvements with respect to the paper [30] can be found in the Master Thesis [43]
@@
Step 1: Proof of $N=1$. We define $\Phi^{\operatorname{NEU}^{1}_{n,\varrho}}_{\varepsilon}=((\operatorname{id},\operatorname{I}_{n},\operatorname{id}))$, from where the result follows easily.
@@
We do this to have upper bounds of the approximations with the cost of the power of 2 as a factor.
@@
$\left\lVert\left(\operatorname{SQR}_{\varepsilon}^{i}(A)+2^{-2^{i}}\operatorname{I}_{n}\right)\operatorname{PROD}_{\varepsilon}^{i}(A)-R\left(\Phi_{\frac{\varepsilon}{4n},1}^{\operatorname{STR}_{n,\varrho}}\right)\left(\operatorname{SQR}_{\varepsilon}^{i}(A)+2^{-2^{i}}\operatorname{I}_{n}\middle|\operatorname{PROD}_{\varepsilon}^{i}(A)\right)\right\rVert_{2}\leq\frac{\varepsilon}{4}.$
@@
We first focus on proving (4.4). By Remark 4.1 with $P=\alpha\operatorname{I}_{n}$, we know that the identity